• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 15:44:35 on June 6,2005

    Entry number 145347

    keyword=Shift Summary - Day

    8:00am Shift starts with work on energy spread at high current.

    8:10 Beam goes away

    8:20 We notice that Left Q3 is no longer regulating. We reboot IOC 10.

    Current (ACC) Events

    4340 60 700k
    4341 60 400k

    4342 junk

    4343 junk

    4344 74 400k
    4345 93 300k


    ---
    Back to second pass beam, energy width is high 5.3E-5 by SLI
    spectrometers are showing a good energy spread (2.8E-5 sigma)
    ---
    Current (ACC)
    4350 5 thin carbon
    4351 5 Ta target
    4352 junk
    4353 5 Ta target SLI approx. 5E-5
    4354 5 thick carbon (OTR in)
    4355 5 thick carbon (OTR out)
    4356 20 thin carbon
    4357 5 waterfall (1350 TACH, 3580) High deadtime
    4358 5 waterfall
    4359 5 waterfall [raster on 1x2]
    4360 junk
    4361 5 thin carbon [raster on 2x2] (beam tripped off)
    4362 5 thin carbon [raster on]
    scanner run 2135

    Bullseye Scan
    -------------

    4363 5 thin carbon [raster off] (position 0,0)
    scanner run 2136

    4364 5 thin carbon (position 1,1)
    scanner run 2137

    4365 5 thin carbon (position -1,1)
    scanner run 2138

    4366 5 thin carbon (position -1,-1)
    scanner run 2139

    4367 5 thin carbon (position 1,-1)
    scanner run 2140

    4368 5 thin carbon (position 0,0)
    scanner run 2141

    ---
    finished, time for access and pass change!

    sieves are going out and the left arm will be put in positive polarity